Green Caterpillar Garden Supplies Inc. is analyzing a project that requires an initial investment of $500,000. The project’s expected cash flows are:

Year Cash Flow Year 1 $275,000 Year 2 –125,000 Year 3 450,000 Year 4 450,000

Green Caterpillar Garden Supplies Inc.’s WACC is 8%, and the project has the same risk as the firm’s average project.

Calculate this project’s modified internal rate of return (MIRR):

19.52%

20.55%

18.50%

22.61%

If Green Caterpillar Garden Supplies Inc.’s managers select projects based on the MIRR criterion, they should ___ this independent project.
Answer #1

For calculating MIRR,

All negative cash flows are discounted to Year 0 at WACC

-500,000 - 125,000/(1+8%)^2 = -$ 607,167.35

All positive cash flows are reinvested to year 4

275,000*(1+8%)^3 + 450,000*(1+8%) + 450,000 = $ 1,282,420.8

607,167.35*(1+MIRR)^4 = 1,282,420.8

MIRR = (1,282,420.8 / 607,167.35)^1/4 -1 = 20.55%

Based on MIRR criterion, as MIRR > Cost of capital , the project should be accepted.